OpenStudy (anonymous):

Determine whether the series is absolutely convergent, conditionally convergent or divergent. Sum (from n=1 to infinity) of [ 2*4*6....(2n)] / n! ]

OpenStudy (turingtest):

finding a representation for [ 2*4*6....(2n)] in terms of n! would be my first move

OpenStudy (anonymous):

lim n-> infinity | | [ 2*4*6... (2n) (2(n+1)) ] / (n+1)! * [ (n!) / (2*4*6 ... 2n) ] | Is this correct?

OpenStudy (anonymous):

Im not sure if I'm supposed to but 2*4*6... (2(n+1)) or 2*4*6... (2n) (2(n+1)) for the numerator of the an+1

OpenStudy (anonymous):

to put*

OpenStudy (turingtest):

2*4*6... (2n) (2(n+1)) is right, but you can't do much with (n+1)! in the denom with the series written this way

OpenStudy (anonymous):

the (n+1)! = (n+1)(n)! right?

OpenStudy (turingtest):

yep, I guess you can deal with this without rewriting the numerator

OpenStudy (anonymous):

and the n! will cancel out?

OpenStudy (turingtest):

right

OpenStudy (anonymous):

so i'll be left with | 2n+2 / n+1 |

OpenStudy (anonymous):

and lim-> infinity of this = 2+0/ 1+0 = 2 > 1 divergent?

OpenStudy (turingtest):

right :)

OpenStudy (anonymous):

alright thank you but how do i determine the second part to prove that it's divergent or conditionally convergent?

OpenStudy (turingtest):

oh I'm rusty on that, but it seems that my earlier idea would show that this series is just divergent rewrite [2*4*6*...*(2n)]=2^n(n!)

OpenStudy (anonymous):

oh okay, thank you for your help!! :)

OpenStudy (turingtest):

if a series is divergent by the ratio test, it is simply divergent I think welcome!

OpenStudy (anonymous):

oh just to note, if its divergent, the non-absolute value has to be divergent in order to be fully divergent. If the non-absolute value is convergent then it's conditionally convergent. :)

OpenStudy (anonymous):

So I guess the non-absolute value of this question is the same.. so it'll be divergent overall. :o hope this is right

OpenStudy (anonymous):

like without the ratio test and maybe manually. Not too sure

OpenStudy (turingtest):

I think you are right since there is not even any sign to consider here but as I said earlier, if you rewrite the numerator as 2^n(n!) then you realize this is just\[\sum_n^{\infty}2^n\] which is clearly divergent